OpenStudy (anonymous):

A canoe can go 24 km downstream in 3 hours. The return trip takes 4 hours. What is the speed of the current? You will earn points for writing the correct system of equations for this word problem and points for solving it correctly and answering the question in a complete sentence.

OpenStudy (victoriasushchik):

So the rate downstream is 8 km/hr and the rate upstream is 6 km/hr (just divide to get these). Let r = rate of canoe in still water and c = rate of current. Thus r + c = 8 r – c = 6 Now add 2r = 14, r = 7. But we want c = 8 – r = 8 – 7 = 1 km/hr The rate of the current is 1 km/hr, ta dah!!

OpenStudy (victoriasushchik):

di tis hewp?

OpenStudy (anonymous):

yessssssssssssss

OpenStudy (victoriasushchik):

you welcome.............